Ch 6.1 What is the formula for calculating an RBRVS charge using the Medicare Physician Fee Schedule?

^
Answer #1

The formula for calculating a RBRVS charge:

Work RVU \(\times\) Work GPCI \(=W\)

Practice Expense RVU \(\times\) Practice Expense \(\mathrm{GPCl}=P E\)

Malpractice RVU \(\times\) Malpractice \(\mathrm{GPCl}=M\)

Conversion Factor \(=C F\)

\((W+P E+M) \times C F=\) Payment

answered by: T.T